how many solutions will this linear system have y=3x+8 and y=3x+2
I have no idea how to solve this so coule you please go step by step in sloving that way i can understand it and be able to solve more problems like this one.
Thanks

Found 2 solutions by Alan3354, Theo:
Answer by Alan3354(69443)

how many solutions will this linear system have y=3x+8 and y=3x+2
-----------
It has none. y can't equal both 3x+8 and 3x+2. If you graph these, you get parallel lines.

Answer by Theo(13342)

equations are:
y = 3x + 8
and
y = 3x + 2
-----
to convert these to standard for for simultaneous equations, you would put the x and y on the left side of the equation and the constant on the right side of the equation.
-----
your equations become:
y - 3x = 8
anbd
y - 3x = 2
-----
if there is a solution, you will have one solution for x that will satisfy both equations. because you have 2 equations in 2 unknowns.
-----
this assumes you need to solve both equations simultaneous meaning that the same solution for x applies to both equations.
-----
immediately you have a problem because you have:
y - 3x = 8
y - 3x = 2
-----
this is because y - 3x can't be equal to 8 and 2 at the same time.
-----
if you tried to solve these equations by subtracting equation 2 from equation 1 you would have gotten the same result.
y - 3x = 8 = equation 1
y - 3x = 2 = equation 2
-----
subtract equation 2 from equation 1 to get:
0 = 6
since this is not a true statement, there is no solution.
-----